#1 A ballpoint pen manufacturer had the following information:

Plastic tubes: top and tip                                 .12 per unit

Ink                                                                       .02 per unit

Direct Labor                                                       .02 per unit

Selling Price                                                      .40 per unit

Advertising                                                        $80,000

Managerial and secretarial salaries               $200,000

Salespeople's commissions Factory                10 per cent of selling price

Factory overhead                                               $120,000

Total available ballpoint pen market is 10 million pens (near this selling price).

Calculate:

a. unit contribution

b. break-even volume in units

c. share of total market to break even

d. total profit for the company if three million pens are sold

e. volume in units required to generate $500,000 profit.

____________________________________________________________________

#2 Using the following information, calculate unit contribution:

a. Advertising = $50,000

b. Break-Even Sales Revenue = $450,000

c. Salaries = $40,000

d. Selling Price = $3 per unit

e. Overhead = $60,000
Answer #1

Answer #1

a. unit contribution $0.20
b. break-even volume in units 2000000
c. share of total market to break even 20%
d. total profit for the company if three million pens are sold $                                                   200,000
e. volume in units required to generate $500,000 profit. 4500000
Detailed working as following
Calculatio of variable product cost and total variable cost'
Plastic tubes: top and tip                                 .12 per unit $0.12
Ink                                                                       .02 per unit $0.02
Direct Labor                                                       .02 per unit $0.02
Variable Product cost $0.16
Variable selling expense = 10% of 0.40 $0.04
Total variable cost per unit $0.20
Fixed Cost = $80000+200000+120000 $                                                   400,000
a. Calculation of Contribution margin
Selling Price per pen $0.40
Less: Variable cost per pen $0.20
Contribution margin $0.20
b. Break even point units =Fixed cost/Contribution per unit
=400000/0.20
2000000
c. Market share of break even =2000000/10000000
20%
d. Income Statement
Sales =3000000*.40 $                                               1,200,000
Less: Variable expenses = 20000 x 49 $                                                   600,000
           Contribution $                                                   600,000
Less: Fixed expense = 26800+12000 $                                                   400,000
Net income $                                                   200,000
e. Volume to generate $500000 profit
Volume to generate $500000 profit =Fixed cost+500000/Contribution per unit
=900000/.20
4500000

Answer #2

contribution per unit = $1

Calculated as

Break even units = 450000/3 = 150000 units

fixed cost = $50000+40000+60000 = $150000

Variable cost per unit = (450000-150000)/150000 = $2

Contribution per unit = Selling price - variable cost per unit = $3-$2 = $1